OpenStudy (anonymous):

Can anyone solve this |x+2|=7

hartnn (hartnn):

when |a| = b, then either a= b or a = -b so what can you say, when |x+2| = 7 ?

OpenStudy (anonymous):

so it would be just x+2=7

hartnn (hartnn):

it would be either x+2 = 7 or x+2 = -7 can you solve those 2 equations ?

OpenStudy (anonymous):

yes thankyou

hartnn (hartnn):

welcome ^_^ if you want me to verify your answers, tell me what u get :)